# non-asymptotic estimates for markov transition matrices with rigorous error bounds

### v1

@@ 4.2. Dimension-free expectation bound
$\mathbb{E}\left[\left\|\bm{R}_{n}-\bm{D}_{\mu}\bm{P}\right\|_{\mathrm{F}}^{2}\middle|u_{1}\sim\nu\right]$ $=\mathbb{E}\left[\left\|\bm{R}_{n}-\bm{D}_{\mu}\bm{P}\right\|_{\mathrm{F}}^{2}\middle|\bm{u}_{1}\sim\nu_{2}\right]$ (4.14)
@@
where $\bm{\Pi}$ is the projection onto $\bm{1}_{d}$ defined in Subsection 2.1. By (2.6), we know that $\bm{P}^{k}-\bm{\Pi}=(\bm{P}-\bm{\Pi})^{k}=(\bm{P}-\bm{\Pi})^{k}(\bm{\mathbf{I}}-\bm{\Pi})$ holds for any $k\in\mathbb{N}_{+}$. As a result, for any integer $k\geq 2$, we have
@@
$\frac{2}{n^{2}}\left\|\sum\limits_{k=1}^{n-1}(n-k)\left(\bm{P}_{2}^{k}-\bm{\Pi}_{2}\right)\right\|_{\mu_{2}}$ $=\frac{2}{n^{2}}\left\|\bm{S}\left((n-1)\bm{\mathbf{I}}+\sum\limits_{k=2}^{n-1}(n-k)\left(\bm{P}-\bm{\Pi}\right)^{k-1}\right)(\bm{\mathbf{I}}-\bm{\Pi})\bm{T}\right\|_{\mu_{2}}$ (4.20)
@@
By the irreducibility of $\bm{P}$, we have $\operatorname{rank}(\bm{\mathbf{I}}-\bm{P})=d-1$. We also know that for any $h\in L_{2,\mu}$, $\operatorname{\mathbb{E}}_{\mu}[(\bm{\mathbf{I}}-\bm{P})h]=0$, and that $(\bm{\mathbf{I}}-\bm{P})\bm{1}=0$. Thus, $\bm{\mathbf{I}}-\bm{P}$ is a bijection from $L_{2,\mu}^{0}$ to $L_{2,\mu}^{0}$. Hence, for any $h\in L_{2,\mu}^{0}$, there exists a function $g\in L_{2,\mu}^{0}$ such that $(\bm{\mathbf{I}}-\bm{P})g=h$. For such $g$, by the definition (2.8) of the symmetric spectral gap $\eta_{s}(\bm{P})$ , we have
$\|g\|_{\mu}^{2}\leq\frac{\left\langle g,(\bm{\mathbf{I}}-\bm{P})g\right\rangle_{\mu}}{\eta_{s}(\bm{P})}=\frac{\left\langle g,h\right\rangle_{\mu}}{\eta_{s}(\bm{P})}\leq\frac{\|g\|_{\mu}\|h\|_{\mu}}{\eta_{s}(\bm{P})},$ (4.23)
which implies $\|g\|_{\mu}\leq\|h\|_{\mu}/\eta_{s}(\bm{P})$. Therefore,
$\left\|\varphi(\bm{P})h\right\|_{\mu}=\left\|\varphi(\bm{P})(\bm{\mathbf{I}}-\bm{P})g\right\|_{\mu}\leq\left\|\varphi(\bm{P})(\bm{\mathbf{I}}-\bm{P})\right\|_{\mu}\left\|g\right\|_{\mu}\leq\frac{\left\|\varphi(\bm{P})(\bm{\mathbf{I}}-\bm{P})\right\|_{\mu}}{\eta_{s}(\bm{P})}\|h\|_{\mu}.$ (4.24)
@@
$\left\|\varphi(\bm{P})h\right\|_{\mu}\leq\frac{4}{n\eta_{s}(\bm{P})}\left\|h\right\|_{\mu},\quad\text{for any}\,\ h\in L_{2,\mu}^{0},$ (4.26)
which implies $\|\varphi(\bm{P}-\bm{\Pi})(\bm{\mathbf{I}}-\bm{\Pi})\|_{\mu}\leq 4/(n\eta_{s}(\bm{P})$. As a result,
$\mathbb{E}\left[\left\|\bm{R}_{n}-\bm{D}_{\mu}\bm{P}\right\|_{\mathrm{F}}^{2}\right]\leq\left(\frac{1}{n}+\left\|\varphi(\bm{P}-\bm{\Pi})(\bm{\mathbf{I}}-\bm{\Pi})\right\|_{\mu}\right)\left(\sum\limits_{\bm{v}\in\Omega^{2}}\left\|\bar{\mathbb{1}}_{\bm{v}}\right\|^{2}_{\mu_{2}}\right)\leq\frac{1}{n}+\frac{4}{n\eta_{s}(\bm{P})}.$ (4.27)
@@
$\left\|\varphi\left(\bm{P}-\bm{\Pi}\right)\right\|_{\mu}$ $=\left\|\varphi\left(\bm{D}_{\mu}^{1/2}(\bm{P}-\bm{\Pi})\bm{D}_{\mu}^{-1/2}\right)\right\|$ (4.31)
@@
Compared with (4.27), the error bound in (4.30) has a better constant. Therefore, (4.30) could be sharper than (4.27) when $\eta_{a}(\bm{P})\geq\eta_{s}(\bm{P})/2$. However, we cannot guarantee that $\bm{P}$ admits a nonzero absolute spectral gap $\eta_{a}(\bm{P})$. In contrast, $\eta_{s}(\bm{P})>0$ holds for any irreducible transition matrix $\bm{P}$. From this point of view, (4.27) is a more robust result.
@@
(4.27) can be view as a matrix version of [4, Theorem 1.2]. In fact, one can apply [4, Theorem 1.2] to $\{\bm{u}_{k}\}_{k\geq 1}$ to obtain an error bound that involves the symmetric spectral gap of $\bm{P_{2}}$. As for our results, we utilize the special structure of $\bm{P}_{2}$ in the proof so that we can use the spectral properties of $\bm{P}$ directly to obtain a sharper error bound.
@@
$\mathbb{E}\left[\left\|\bm{H}_{n}-\bm{D}_{\mu}\bm{P}\right\|_{\mathrm{F}}^{2}\middle|\tilde{\bm{u}}_{1}\sim{\tilde{\nu}_{2}}\right]$ $\leq\|\tilde{\nu}_{2}/\tilde{\mu}_{2}\|_{\infty}\mathbb{E}\left[\left\|\bm{H}_{n}-\bm{D}_{\mu}\bm{P}\right\|_{\mathrm{F}}^{2}\middle|\tilde{\bm{u}}_{1}\sim{\tilde{\mu}_{2}}\right]$ (4.33)
@@
$\left\|\tilde{\varphi}\left(\tilde{\bm{P}}_{2}-\tilde{\bm{\Pi}}_{2}\right)\right\|_{\tilde{\mu}_{2}}$ $\leq\tilde{\varphi}\left(\left\|\tilde{\bm{P}}_{2}-\tilde{\bm{\Pi}}_{2}\right\|_{\tilde{\mu}_{2}}\right)=\tilde{\varphi}\left(\lambda_{a}\left(\tilde{\bm{P}}_{2}\right)\right)$ (4.38)

### v2

@@ 4.2. Dimension-free expectation bound
$\mathbb{E}\left[\left\|\bm{R}_{n}-\bm{D}_{\mu}\bm{P}\right\|_{\mathrm{F}}^{2}\middle|u_{1}\sim\nu\right]$ $=\mathbb{E}\left[\left\|\bm{R}_{n}-\bm{D}_{\mu}\bm{P}\right\|_{\mathrm{F}}^{2}\middle|\bm{u}_{1}\sim\nu_{2}\right]$
@@
where $\bm{\Pi}$ is the projection onto $\bm{1}_{d}$ defined in Subsection 2.1. One can verify that $\bm{P}^{k}-\bm{\Pi}=(\bm{P}-\bm{\Pi})^{k}=(\bm{P}-\bm{\Pi})^{k}(\bm{\mathbf{I}}-\bm{\Pi})$ holds for any $k\in\mathbb{N}_{+}$. As a result, for any integer $k\geq 2$, we have
@@
$\frac{2}{n^{2}}\left\|\sum\limits_{k=1}^{n-1}(n-k)\left(\bm{P}_{2}^{k}-\bm{\Pi}_{2}\right)\right\|_{\mu_{2}}$ $=\frac{2}{n^{2}}\left\|\bm{S}\left((n-1)\bm{\mathbf{I}}+\sum\limits_{k=2}^{n-1}(n-k)\left(\bm{P}-\bm{\Pi}\right)^{k-1}\right)(\bm{\mathbf{I}}-\bm{\Pi})\bm{T}\right\|_{\mu_{2}}$
@@
By the irreducibility of $\bm{P}$, we have $\operatorname{rank}(\bm{\mathbf{I}}-\bm{P})=d-1$. We also know that for any $h\in L_{2,\mu}$, $\operatorname{\mathbb{E}}_{\mu}[(\bm{\mathbf{I}}-\bm{P})h]=0$, and that $(\bm{\mathbf{I}}-\bm{P})\bm{1}=0$. Thus, $\bm{\mathbf{I}}-\bm{P}$ is a bijection from $L_{2,\mu}^{0}$ to $L_{2,\mu}^{0}$. Hence, for any $h\in L_{2,\mu}^{0}$, there exists a function $g\in L_{2,\mu}^{0}$ such that $(\bm{\mathbf{I}}-\bm{P})g=h$. For such $g$, by the definition of the IP gap $\eta_{p}(\bm{P})$, we have $\|g\|_{\mu}\leq\|h\|_{\mu}/\eta_{P}(\bm{P})$. Therefore,
$\left\|\varphi(\bm{P})h\right\|_{\mu}=\left\|\varphi(\bm{P})(\bm{\mathbf{I}}-\bm{P})g\right\|_{\mu}\leq\left\|\varphi(\bm{P})(\bm{\mathbf{I}}-\bm{P})\right\|_{\mu}\left\|g\right\|_{\mu}\leq\frac{\left\|\varphi(\bm{P})(\bm{\mathbf{I}}-\bm{P})\right\|_{\mu}}{\eta_{p}(\bm{P})}\|h\|_{\mu}.$ (4.21)
@@
$\left\|\varphi(\bm{P})h\right\|_{\mu}\leq\frac{4}{n\eta_{p}(\bm{P})}\left\|h\right\|_{\mu},\quad\text{for any}\,\ h\in L_{2,\mu}^{0},$ (4.23)
which implies $\|\varphi(\bm{P}-\bm{\Pi})(\bm{\mathbf{I}}-\bm{\Pi})\|_{\mu}\leq 4/(n\eta_{p}(\bm{P}))$. As a result,
$\mathbb{E}\left[\left\|\bm{R}_{n}-\bm{D}_{\mu}\bm{P}\right\|_{\mathrm{F}}^{2}\right]\leq\left(\frac{1}{n}+\left\|\varphi(\bm{P}-\bm{\Pi})(\bm{\mathbf{I}}-\bm{\Pi})\right\|_{\mu}\right)\left(\sum\limits_{\bm{v}\in\Omega^{2}}\left\|\bar{\mathbb{1}}_{\bm{v}}\right\|^{2}_{\mu_{2}}\right)\leq\frac{1}{n}+\frac{4}{n\eta_{p}(\bm{P})}.$ (4.24)
@@
$\left\|\varphi\left(\bm{P}-\bm{\Pi}\right)\right\|_{\mu}$ $=\left\|\varphi\left(\bm{D}_{\mu}^{1/2}(\bm{P}-\bm{\Pi})\bm{D}_{\mu}^{-1/2}\right)\right\|$
@@
Whether the error depends on the dimension shall rely on the choice of the error metric. The reason we can obtain dimension free expectation bound is that we measure the error using the Frobenius norm (i.e. $L_{2,u})$, which is a natural choice for matrices. If one instead uses the TV norm (i.e. $L_{1,u}$) to measure the error, a dependence on the dimension typically arises. This is why in [23], the error bound depends on the quantity $\|\bm{D}_{\mu}\bm{P}\|_{1/2}=(\sum_{u,v\in\Omega}(\mu(u)p(u,v))^{1/2})^{2}$, which may depend on the dimension $d$ and can be bounded by $d^{2}$.
To better illustrate this point, consider estimating a discrete distribution on a state space of size $d$ by the empirical distribution of a large amount of i.i.d samples. Then one can see that the $L_{2}$ error between the empirical distribution and the ground truth can be bounded independently of $d$. In contrast, the error in the TV norm scales like $d^{1/2}$ in the worst case.
@@
It is worth noting that Theorem 2.1 can be extended to Markov chains on a countable state space, provided that the transition operator $\bm{P}$ admits a nonzero IP gap $\eta_{p}(\bm{P})>0$. The only essential modification appears in the proof of Theorem 2.1(3) where we need to show that the operator $\bm{I}-\bm{P}$ is invertible on $L_{2,\mu}^{0}$. In the finite-dimensional case, this is straightforward, since $\bm{I}-\bm{P}$ is one to one. In the countable (infinite-dimensional) case, we can prove this by contradiction. Suppose $\bm{I}-\bm{P}$ is not invertible on $L_{2,\mu}^{0}$. Then there exists a nonzero function $g\in L_{2,\mu}^{0}$ such that for all $h\in L_{2,\mu}^{0}$, $\langle g,(\bm{\mathbf{I}}-\bm{P})h\rangle_{\mu}=0$. In particular, by taking $h=g$, we deduce that $\|g\|_{\mu}^{2}=\langle g,\bm{P}g\rangle_{\mu}$. On the other hand, we have
$\langle g,\bm{P}g\rangle_{\mu}\leq\|g\|_{\mu}\|\bm{P}g\|_{\mu}\leq\|g\|_{\mu}^{2},$
so equality must hold throughout, which implies $(\bm{\mathbf{I}}-\bm{P})g=0$. This contradicts with the fact that $\eta_{p}(\bm{P})>0$.
Remark 4.5.
The estimate in (4.24) can be view as a matrix version of [5, Theorem 1.2].In fact, one can apply [5, Theorem 1.2] to $\{\bm{u}_{k}\}_{k\geq 1}$ to obtain that
$\mathbb{E}\left[\left\|\bm{R}_{n}-\bm{D}_{\mu}\bm{P}\right\|_{\mathrm{F}}^{2}\right]\lesssim\frac{1}{n\eta_{p}(\bm{P}_{2})}\lesssim\frac{1}{n\eta_{p}(\bm{P})}.$
The second inequality above owes to $\eta_{p}(\bm{P}_{2})\geq\eta_{p}(\bm{P})/(1+\eta_{p}(\bm{P}))$ in Theorem 3.3. In contrast, the proof we present above is more straightforward in the sense that it controls the expected error in the Frobenius norm directly by $1/(n\eta_{p}(\bm{P}))$ without explicitly establishing the relation between $\eta_{p}(\bm{P}_{2})$ and $\eta_{p}(\bm{P})$.
Remark 4.6.
Compared with (4.24), the error bound in (4.27) has a better constant. Therefore, (4.27) could be sharper than (4.24) when $\eta_{a}(\bm{P})\geq\eta_{p}(\bm{P})/2$. However, we cannot guarantee that $\bm{P}$ admits a nonzero absolute spectral gap $\eta_{a}(\bm{P})$. In contrast, $\eta_{p}(\bm{P})>0$ holds for any irreducible transition matrix $\bm{P}$. From this point of view, (4.24) is a more robust result.
@@
$\mathbb{E}\left[\left\|\bm{H}_{n}-\bm{D}_{\mu}\bm{P}\right\|_{\mathrm{F}}^{2}\middle|\tilde{\bm{u}}_{1}\sim{\tilde{\nu}_{2}}\right]$ $\leq\|\tilde{\nu}_{2}/\tilde{\mu}_{2}\|_{\infty}\mathbb{E}\left[\left\|\bm{H}_{n}-\bm{D}_{\mu}\bm{P}\right\|_{\mathrm{F}}^{2}\middle|\tilde{\bm{u}}_{1}\sim{\tilde{\mu}_{2}}\right]$
@@
$\left\|\tilde{\varphi}\left(\tilde{\bm{P}}_{2}-\tilde{\bm{\Pi}}_{2}\right)\right\|_{\tilde{\mu}_{2}}$ $\leq\tilde{\varphi}\left(\left\|\tilde{\bm{P}}_{2}-\tilde{\bm{\Pi}}_{2}\right\|_{\tilde{\mu}_{2}}\right)=\tilde{\varphi}\left(\lambda_{a}\left(\tilde{\bm{P}}_{2}\right)\right)$
@@
Appendix A Comparison of different spectral gaps
This section provides a brief comparison of different spectral gaps introduced in Section 2. Since all of these results already appear in [5, 10], we present them here solely for the convenience of the reader. We first show that $\eta_{p}(\bm{P})\geq\eta_{s}(\bm{P})\geq\eta_{a}(\bm{P})$ and $\eta_{p}(\bm{P})\geq\eta_{ps}(\bm{P})/2$. As a direct consequence, our main results also holds if $\bm{P}$ admits a nonzero absolute spectral gap, a nonzero pseudo spectral gap,or a nonzero symmetric spectral gap.
Lemma A.1.
It holds that $\eta_{p}(\bm{P})\geq\eta_{s}(\bm{P})\geq\eta_{a}(\bm{P})$ and $\eta_{p}(\bm{P})\geq\eta_{ps}(\bm{P})/2$ for any irreducible transition $\bm{P}$ with an invariant distribution $\mu$.
Proof.
For any $h\in L_{2,\mu}^{0}$, by the Cauchy–Schwarz inequality,
$\frac{\langle h,(\bm{\mathbf{I}}-\bm{P})h\rangle_{\mu}}{\|h\|_{\mu}^{2}}\leq\frac{\|(\bm{\mathbf{I}}-\bm{P})h\|_{\mu}}{\|h\|_{\mu}}.$
Taking the supremum over $h\in L_{2,\mu}^{0}$, we deduce $\eta_{p}(\bm{P})\geq\eta_{s}(\bm{P})$. Similarly, since
$\frac{\langle h,\bm{P}h\rangle_{\mu}}{\|h\|_{\mu}^{2}}\leq\frac{\|\bm{P}h\|_{\mu}}{\|h\|_{\mu}}$
holds for any $h\in L_{2,\mu}^{0}$, we deduce $\lambda_{s}(\bm{P})\leq\lambda_{a}(\bm{P})$, and therefore $\eta_{s}(\bm{P})\geq\eta_{a}(\bm{P})$.
To prove $\eta_{p}(\bm{P})\geq\eta_{ps}(\bm{P})/2$, we first note that for any $k\in\mathbb{N}_{+}$,
$\left\|\left(\bm{P}^{k-1}-\bm{P}^{k}\right)h\right\|_{\mu}=\left\|\bm{P}^{k-1}(\bm{\mathbf{I}}-\bm{P})h\right\|_{\mu}\leq\|(\bm{\mathbf{I}}-\bm{P})h\|_{\mu}.$
Then, for any $h\in L_{2,\mu}^{0}$ and any $k\in\mathbb{N}_{+}$, we can compute by triangle inequality and telescoping that
$1-\frac{\left\|\bm{P}^{k}h\right\|_{\mu}}{\|h\|_{\mu}}\leq\frac{\left\|\left(\bm{\mathbf{I}}-\bm{P}^{k}\right)h\right\|_{\mu}}{\|h\|_{\mu}}\leq\frac{1}{\|h\|_{\mu}}\sum\limits_{i=1}^{k}\left\|\left(\bm{P}^{i-1}-\bm{P}^{i}\right)h\right\|_{\mu}\leq k\frac{\|(\bm{\mathbf{I}}-\bm{P})h\|_{\mu}}{\|h\|_{\mu}}.$
It follows by taking the infimum over all $h\in L_{2,\mu}$ that $1-\lambda_{a}(\bm{P}^{k})\leq k\eta_{p}(\bm{P})$, and thus
$\eta_{ps}(\bm{P})=\sup_{k\in\mathbb{N}_{+}}\frac{1-\lambda_{a}^{2}\left(\bm{P}^{k}\right)}{k}\leq\sup_{k\in\mathbb{N}_{+}}\frac{2\left(1-\lambda_{a}\left(\bm{P}^{k}\right)\right)}{k}\leq 2\eta_{p}(\bm{P}).$
We have used above the fact that $\lambda_{a}(\bm{P}^{k})\leq 1$ (2.15). ∎
Recall that the irreducible transition matrix $\bm{P}_{2}$ constructed in Section 3 is already an example whose IP gap is positive while whose absolute spectral gap is zero. Below, we provide two more examples demonstrating the superiority of the IP gap.
Example A.2.
Chatterjee provided in [5] an example demonstrating that the ratio $\eta_{a}{\bm{P}}/\eta_{p}(\bm{P})$ can be arbitrarily small. Consider the one dimensional random walk on the periodic grid $\mathbb{Z}/n\mathbb{Z}$ where the probability of not moving and that of moving one step to the right are both $1/2$. It is then not hard to check that the IP gap of this Markov chain is of order $1/n$, while the symmetric spectral gap and the absolute spectral are both of order $1/n^{2}$.
Example A.3 (Perturbation of a periodic transition matrix).
Consider a parameterized family of transition matrices,
$\bm{P}_{0}=\begin{bmatrix}0&1&0\\ 0&0&1\\ 1&0&0\\ \end{bmatrix},\quad\text{and }\quad\bm{P}_{\varepsilon}=\varepsilon\bm{\mathbf{I}}+(1-\varepsilon)\bm{P}_{0},\quad\varepsilon\in[0,1].$ (A.1)
Suppose $\varepsilon\ll 1$. One can easily show by a straightforward calculation and the spectral perturbation theory that $\eta_{p}(\bm{P}_{\varepsilon})=(1-\varepsilon)\sqrt{3}$, $\eta_{s}(\bm{P}_{\varepsilon})=3(1-\varepsilon)/2$, $\eta_{ps}(\bm{P}_{\varepsilon})=\mathcal{O}(\varepsilon)$, $\eta_{a}(\bm{P}_{\varepsilon})=\mathcal{O}(\varepsilon)$. In particular, $\eta_{p}(\bm{P}_{0})=\sqrt{3}$, $\eta_{s}(\bm{P}_{0})=3/2$ and $\eta_{a}(\bm{P}_{0})=\eta_{ps}(\bm{P}_{0})=0$.